OpenStudy (anonymous):

A spherical balloon is inflated so that it's volume is increasing at A rate of 2 m^3/sec. At what rate is the balloon's radius increasing when the radius is 7 m?

myininaya (myininaya):

V=4/3 *pi *r^3 is volume of sphere V'=4/3*pi*3r^2*r' is the rate of the volume (we are given V'=2 m^3/sec) we want to find r' when r=7. 2=4/3*pi*3*(7^2)*r' 2=4*pi*49*r' 2=196pi*r' 2/(196pi)=r'

myininaya (myininaya):

r'=1/(98pi)

OpenStudy (anonymous):

Thank you! Could you please help me answer the newton's method question I posted earlier?